OpenStudy (anonymous):

A projectile is fired horizontally from a gun that is 70.0 m above flat ground, emerging from the gun with a speed of 250 m/s. (a) How long does the projectile remain in the air? s (b) At what horizontal distance from the firing point does it strike the ground? m (c) What is the magnitude of the vertical component of its velocity as it strikes the ground?

OpenStudy (surry99):

Start by drawing a labeled diagram

OpenStudy (abhisar):

#a Since the projectile is thrown horizontally, vertical component of the velocity will be zero. Thus you can calculate the time of flight (t) by using \(\sf s=ut+\frac {1}{2}at^2\) Take, s=70m u=0 a=9.8

OpenStudy (abhisar):

Once you find t in the above question, use it to find horizontal distance (s) by using the formula \(\sf s=ut+\frac{1}{2}at^2\) take, u=250m/s t= as calculated in the above problem a= 0 (because there is no acceleration in the horizontal direction)

OpenStudy (abhisar):

#c Finally calculate the final velocity (v) by using the equation v=u+at Take, u=0 t= as calculated in the first problem a=9.8

OpenStudy (anonymous):

thanks that was very helpful for many questions not just this one